Dr Paul Coddington

Northeast Parallel Architectures Center 111 College Place, 3-228 CST Syracuse University Syracuse, NY 13244

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

January 10, 1994

Nelson Simons Power Systems Engineering Department General Electric Company 1 River Road, Bldg. 2, Room 544 Schenectady, NY 12345

Dear Nelson,

Thank you for the information on MAPS. I will review it with some students and a research scientist who I hope will be able to work on this project. I have signed and enclosed the requested non-disclosures.

I will try to get back to you by the end of the week with some possible times for us to visit you during the week of January 24th. During that visit I would also like to discuss in more detail creating a proposal to submit to the NYS Energy Research and Development Authority. I will work with the Office of Sponsored Programs at Syracuse University on drafting a proposal.

Yours sincerely,

Paul Coddington